OpenStudy (anonymous):

what is the vertex for y = x2 - 2x - 3

OpenStudy (fanduekisses):

vertex: \[-\frac{ b }{ 2a }\] where, y = ax^2 + bx + c.

OpenStudy (fanduekisses):

b=-2 and a=1, so \[-\frac{ -2 }{ 2*1 }\]

OpenStudy (fanduekisses):

actually, -b/2a just gives the x-coordinate of the vertex

OpenStudy (fanduekisses):

to find the y-coordinate just substitute the x which is 1 so y=1^2-2*1-3=-4

OpenStudy (fanduekisses):

so vertex: (1,-4)